OpenStudy (anonymous):

State whether the slope of a line passing through (2, -3) and (1, 8) is positive, negative, zero or undefined.

OpenStudy (anonymous):

i put negative was i right?

OpenStudy (alexandervonhumboldt2):

use slope formula

OpenStudy (alexandervonhumboldt2):

\[\frac{ y_1-y_2 }{ x_1-x_2 }\]

OpenStudy (alexandervonhumboldt2):

\[\frac{ -3-8 }{ 2-1 }=negative~~slope\]

OpenStudy (anonymous):

= -11

OpenStudy (alexandervonhumboldt2):

yeaeh

OpenStudy (anonymous):

so it is negative

OpenStudy (alexandervonhumboldt2):

yeah